# Potential Contrast: Properties, Equivalences, and Generalization to Multiple Classes

### v1

@@ V Application
Figure 3: From top to bottom. 1. A grayscale image showing an excerpt from GB-Cgc 667/760, page 19, with foreground notation (red), bleedthrough (green), and background (blue) labeled. 2. A three-class NPC result with foreground in black, bleedthrough in gray, and background in white. The three-class NPC value is 0.965. The two-class NPC between foregreound notation and background is nearly 1, the two-class NPC between foreground and bleedthrough is 0.996, and the two-class NPC between bleedthrough and background is 0.934. 3. The three-class NPC segmentation showing only foreground notation. Original imaging by DIAMM (%4.␣The␣three-class␣NPC␣result␣showing␣only␣bleedthrough.diamm.ac.uk). Image use by kind permission of the Master and Fellows of Gonville and Caius College, Cambridge.
To illustrate the utility of multi-class NPC, we apply it to photographed page from the Caius Choirbook (GB-CGC 667/760), a music manuscript dated to the late 1520s and currently at Gonville and Caius College, Cambridge [11]. Like many historical manuscripts, this choirbook exhibits substantial bleedthrough, showing ink from the back of the page as well as the front. We apply NPC with three classes: written notation on the open page, bleedthrough, and background. The result, shown in Figure 3, is a strong multi-class NPC value, suggesting that the labeled pixels are easily separated using only their intensity values. This is reflected in the segmentation, where foreground text is easily isolated, allowing us to remove the bleedthrough from the page and visualize the music without undesired noise from the background.
The two-class NPC values also reflect what we expect: the foregrorund is easy to separate from background and from bleedthrough, but separating background from bleedthrough is more difficult. However, for many applications a single value is more useful, accounting for all classes rather than only pairwise comparisons.
@@
Finally, the NPC value also heavily depends on the labeled pixels. While labeling the pixels can take time, usually only a small number of pixels must be labeled in each class. This labeling also allows allows a user or domain expert to adapt the measure to a particular task.
@@ VI Conclusion
Potential contrast is an image contrast and quality measure that relies on task-dependent labels. We introduce a scaled version that is commensurate across image formats and applications, which retains many of the desirable properties of potential contrast.
We prove equalities that support different interpretations of normalized potential contrast and show it is equivalent to the total variation distance on the probability distributions of sampled pixels, allowing generalization to continuous domains. We use another equality to define normalized potential contrast when there are more than two classes.
Such generalizations enable applications to a much broader array of contexts, including where potential contrast has already found success. Multi-class NPC can provide a valuable, interpretable, task-adaptive image quality and contrast measure in cultural heritage and beyond.

### v2

@@ V Application
Figure 3: (Top) A grayscale image showing an excerpt from GB-Cgc 667/760, page 19, with foreground notation (red), bleedthrough (green), and background (blue) labeled. (Middle) A three-class NPC result with foreground in black, bleedthrough in gray, and background in white. The three-class NPC value is 0.965. The two-class NPC between foreground notation and background is nearly 1, the two-class NPC between foreground and bleedthrough is 0.996, and the two-class NPC between bleedthrough and background is 0.934. (Bottom) The three-class NPC segmentation showing only foreground notation. Original imaging by DIAMM (%4.␣The␣three-class␣NPC␣result␣showing␣only␣bleedthrough.diamm.ac.uk). Image use by kind permission of the Master and Fellows of Gonville and Caius College, Cambridge.
To illustrate the utility of multi-class NPC, we apply it to photographed page from the Caius Choirbook (GB-CGC 667/760), a music manuscript dated to the late 1520s and currently at Gonville and Caius College, Cambridge [11]. Like many historical manuscripts, this choirbook exhibits substantial bleedthrough, showing ink from the back of the page as well as the front. We apply NPC with three classes: written notation on the open page, bleedthrough, and background. The result, shown in Figure 3, is a strong multi-class NPC value, suggesting that the labeled pixels are easily separated based only on their intensity values. This is reflected in the segmentation, where foreground text is easily isolated, allowing us to remove the bleedthrough from the page and visualize the music without undesired noise from the background.
The two-class NPC values also reflect what we expect: the foregrorund is easy to separate from background and from bleedthrough, but separating background from bleedthrough is more difficult. However, for many applications a single value is more useful, one which accounts for all classes rather than only pairwise comparisons.
@@
Finally, the NPC value also heavily depends on the labeled pixels. While labeling can take time, usually only a small number of pixels are needed from each class. This labeling also allows a user to adapt the measure to a particular task.
@@ VI Conclusion
Potential contrast is an image contrast and quality measure that relies on task-dependent labels. We introduce a scaled version, normalized potential contrast (NPC), that is commensurate across image formats and applications, which retains many of the desirable properties of potential contrast.
We prove equalities that support different interpretations of NPC and show it is equivalent to the total variation distance on the probability distributions of sampled pixels, allowing generalization to continuous domains. One of the other equivalences allows us to define NPC when there are more than two classes, namely multi-class NPC.
Such generalizations enable applications to a much broader array of contexts, including where potential contrast has already found success. We showcase an example with three classes in a music manuscript exhibiting bleedthrough, demonstrating how multi-class NPC can provide a valuable, interpretable, and task-adaptive image quality and contrast measure in cultural heritage and beyond.